OpenStudy (anonymous):

Evaluate the expression 4a − 6b + c for a = 6, b = 3, c = 8 A. −14 B. 14 C. 11 D. −11

OpenStudy (pokemon23):

should have learned it in daycare

OpenStudy (pokemon23):

4(6)-6(3)+8

OpenStudy (anonymous):

24 - 18 + 8 = 14

OpenStudy (lilg132):

4(6) - 6(3) + 8 = 24 - 18 + 8 = 14

OpenStudy (lilg132):

B 14

OpenStudy (pokemon23):

answer is 14

OpenStudy (anonymous):

LETS BE HOONEST here its like free medals :)